OpenStudy (anonymous):

Can someone explain how in Problem Set 1, Problem 1I-3 how the function equals 1 when n -> infinity. I'm having problem seeing this because LN(1) = zero.

OpenStudy (anonymous):

Just in Case. The problem is Evaluate LIM(n->infinity) n*LN(1+1/n) Hint: Let h = 1/n, and use (d/dx)ln(1+x), x=0, = 1

OpenStudy (anonymous):

It's explained during Part B of Unit 1, starting with Session 16, I think. Note that while ln(1+1/n) tends towards zero, n tends towards infinity, so n*ln(1+1/n) tends towards something like infinity times zero.